I dont understand how to do Unit Rate at all. Could someone help me on giving me step by step instructions how to?
Answer by gonzo(654)

unit rate * time = number of units
the above equation is a necessary starting point.
working from that.....
let unit rate = R.
let number of units = U
let time = T
the algebraic form of the equation is then R*T=U
the unit rate is expressed in units per time
as an example:
let U = 3000
let R = 50 units per hour
let T = total number of hours
putting the values in the formula R*T=U, we get
50*T=3000
solving for T, we get
T=3000/50
T=60 hours
proving the formula is correct, we substitute all values into the original equation.
since R*T=U, and R = 50 units per hour, T = 60 hours, U = 3000 units, we get
50*60=3000 which translates to
50 units per hour * 60 hours produces 3000 units.
since 50*60 = 3000, the equation becomes
3000 = 3000
which proves the formula is correct.